OpenStudy (anonymous):

Ginny's class ordered 3 pizzas. They ate five-twelfths of each pizza. How many pizzas did they eat?

OpenStudy (anonymous):

5/12+5/12+5/12=15/12=5/4=1and1/4

OpenStudy (anonymous):

What Johnny said ...

OpenStudy (anonymous):

5/4=1.25